1.graph the height of a right triangle as a function of the base ,if the hypotenuse is fixed at 10.


2.graph the diameter of a circle as a function of it's circumstances.


3.determine whether the following function is continuous or discontinuous .


f(x)=3x+2/x^2-6


a.at x =-1


b.at x =-2


1
Expert's answer
2022-02-24T17:01:33-0500

1.

"y=\sqrt{100-x^2},\{0<x<10\}"



2.


"y=\dfrac{x}{\pi}, x\ge0"


3.


"x^2-6\not=0"

"x=-\sqrt{6}, x=\sqrt{6}"

a.


"f(-1)=\dfrac{3(-1)+2}{(-1)^2-6}=\dfrac{1}{5}"

"\lim\limits_{x\to -1}\dfrac{3x+2}{x^2-6}=\dfrac{3(-1)+2}{(-1)^2-6}=\dfrac{1}{5}=f(-1)"


The function "f(x)" is continuous at "x=-1."


b.


"f(-2)=\dfrac{3(-2)+2}{(-2)^2-6}=2"

"\lim\limits_{x\to -2}\dfrac{3x+2}{x^2-6}=\dfrac{3(-2)+2}{(-2)^2-6}=2=f(-2)"


The function "f(x)" is continuous at "x=-2."